OpenStudy (anonymous):

i need help in algebra 1 foundations A CA

OpenStudy (anonymous):

Ask away, friend.

OpenStudy (anonymous):

alright hang on

OpenStudy (anonymous):

write an equation in point slope form of the line through the given point and with the given slope. y=3/4 x+5

OpenStudy (anonymous):

5-8 = 3/4 (0-4)

OpenStudy (anonymous):

(-2, 3);m=-2

OpenStudy (anonymous):

Wha?

OpenStudy (anonymous):

i dont know any of this and if i dont do good im not going to pass please help

OpenStudy (anonymous):

do the same for the next one as you did the last one

OpenStudy (anonymous):

(3-1)=-2(-2-(-1))